### v1

@@ 5.4 Prediction performance analysis
Then we compute the average RMSE value tested on four datasets before and after optimization using our approach, illustrating the improvement of performance of reservoirs. The RMSE is computed as
@@
where $n$ is the time steps to predict, $y_{i}$ and $\hat{y}_{i}$ represent the predict value and real value of the $i$-step respectively. Table 2 shows the train and test time on the four datasets  [^3] All the experiments were conducted on a server equipped with dual Intel(R) Xeon(R) CPU E5-2678 v3 processors (2.50GHz, 24 cores, 48 threads) and 62 GB of RAM., and Fig. 8 summarizes our experimental results, which shows the tested results on four datasets using three different initial construction. Here, for the Random initialization method, the initial reservoir digraphs are sparse and hence show weak memory capacities and prediction abilities at first, even if the dataset has obvious topological features. However, our experimental results demonstrate that even highly sparse reservoirs can achieve substantially improved prediction performance through adding the number of rings within their digraph architecture. And for the small-world and scale-free methods, since in some studies they are proved to be effective initialization ways of reservoirs [44, 45, 19], we find that they have already achieve relative good performance of prediction on some datasets. But for these cases, we have also observed that after adding the number of rings, if the original reservoir already possessed great prediction capability (such as the small-world and scale-free method on the Lorenz system and NARMA datasets), the modified results can maintain such good capability, i.e., obtaining similar RMSE or lower RMSE. On the contrary, for reservoirs with initially poor prediction capabilities (such as the Mackey-Glass dataset), the optimization process proves to be even more useful, where the modified reservoirs experience a significant enhancement in their prediction power.
@@
Mackey-Glass 2.1 1.7 3.0 2.7 2.7 2.5
MSO 2.2 1.8 2.7 2.6 2.9 2.6
Lorenz System 2.2 1.8 2.7 2.5 2.6 2.5
NARMA 2.3 1.9 3.0 2.6 2.8 2.8
Moreover, our optimization method yields varying predictive abilities across datasets with distinct topological features. Specifically, the Mackey-Glass dataset demonstrated the most significant improvement in predictive performance. As illustrated by the persistence diagram’s confidence set (Fig. 5), the Mackey-Glass dataset exhibits clear loop features in its embedding space. This aligns perfectly with the conditions under which our method excels, as it leverages these loop structures to enable the optimized reservoir digraph to capture the system’s periodic behaviors effectively. Conversely, the improvement observed with the NARMA dataset was comparatively modest. The NARMA dataset lacks pronounced loop features, which limits our method’s ability to identify strong topological characteristics. Consequently, even after optimization, the performance enhancement remained constrained. The MSO and Lorenz System datasets showed intermediate levels of improvement, falling between the Mackey-Glass and NARMA datasets. While these datasets primarily exhibit noisy topological features, they contain some points with relatively greater persistence compared to the NARMA dataset. Indeed, periodic patterns are discernible in their embedded point clouds. Hence, these two datasets experienced a moderate improvement in the reservoir digraph’s predictive ability after optimization.
Figure 8: Average RMSE tested on four datasets before and after optimizing the reservoir digraph. Here we compared three initialization approaches.
@@ 5.5 Discussion
By combining the topological features of the dataset shown in Fig. 5, we found that, under our experimental conditions, a clear pattern emerged regarding the performance of our method: our method demonstrated relatively good performance for datasets with obvious 1-dimensional topological features, namely loops, in the time-delay embedding space. This aligns with the statement in section 4.1, i.e., the reservoir usually performs well on the datasets with pronounced periodicity. In the context of reservoir computing, these 1-d loops represent recurrent structures that can capture temporal dependencies and dynamic behaviors within the data. To predict such kind of dataset (like the Mackey-Glass dataset), successful modeling requires accurate detection and storage of periodic patterns. Reservoir computing architectures are particularly well-suited for this purpose, as their memory capacity enables effective capture of temporal dependencies. This explains why reservoirs with enhanced MC through our optimization method demonstrate good performance on such datasets. When applied to such datasets, our method effectively interacts with these periodic patterns. These features enable the reservoir to better extract and utilize the sequential information, facilitating more accurate representation and forecasting of the underlying system dynamics. Conversely, for datasets with indistinct topological features, the enhancement brought about by our method to the prediction ability of the reservoir digraph was relatively limited. In these cases, the lack of prominent topological structures, such as significant loops, means that the reservoir has fewer obvious patterns to extract when predicting. Without clear topological features, the method’s impact on improving the model’s performance is restricted.
Hence, all these phenomena can be comprehensively understood as the effectiveness of our method being the result of the joint influence of the structural optimization of the reservoir and the characteristics of the dataset itself. On one hand, our method modifies the internal architecture of the reservoir digraph using GLMY homology. On the other hand, the inherent topological features of the dataset act as a foundation, determining the extent to which the method can optimize the model. Only when the method can effectively identify and utilize the topological features of the dataset, as in the case of datasets with distinct one-dimensional loops, can it achieve significant improvements in the prediction performance of the reservoir digraph. Thus, if there is a clear periodic pattern (1-dimensional topological features) in the given time series data, using our optimization method allows the reservoir to generate stronger predictive ability.
In summary, our study not only reveals the performance characteristics of our method under different dataset conditions but also provides valuable insights for future research on integrating topological analysis with model optimization in complex system modeling.
@@
There are some limitations of our method. First, our approach does not capture or modify boundary triangles and boundary squares, as they do not represent generators of GLMY homology despite being cycles. In dense reservoir digraphs, as the connection density is large, boundary triangles can be numerical, making the number of representative cycles be small. But for a sparse reservoir digraph, the boundary cycles are significantly less than the representative cycles, the majority of cycles can still be transformed into rings, thereby enhancing the reservoir’s prediction capability. Hence, our method is more suitable for optimizing sparse reservoirs. And since many rings can share common edges (which means their corresponding fundamental cycle submatrices will share common columns), the increased number of rings is not proportional to improved orthogonality. It is still worthy studying more efficient and reasonable method for improving orthogonality.
@@
Additionally, it is still difficult to give a complete theoretical explanation for why our method can enhance the effect of reservoir, especially when reservoir is a large-scale random sparse matrix, because its properties are difficult to study. Moreover, the computational efficiency also becomes a concern for dense reservoirs, as computing GLMY homology and identifying representative cycles can be time-intensive. Our future research directions will focus on investigating the theoretical relationship between reservoir topology and computational capability, and developing accelerated algorithms for GLMY homology computations.
@@ 6 Conclusions
In this study, we proposed a novel approach to optimize the structure of the reservoirs by GLMY homology theory. We begin by computing the one-dimensional GLMY homology of the reservoir digraph to obtain the corresponding minimal representative cycles. These cycles are then utilized to strategically modify the digraph structure in a compatible way, specifically increasing the number of rings within the reservoir. Experimental results demonstrate that this topological optimization leads to improved predictive performance across our tested datasets, and the proposed method performs better on datasets with significant 1-dimensional topological features.

### v2

@@ 5.5 Prediction performance analysis
Then we compute the average RMSE and standard deviation tested on five datasets before and after optimization using our approach, illustrating the improvement in the performance of reservoirs. The RMSE is computed as
@@
where $n$ is the time steps to predict, $y_{i}$ and $\hat{y}_{i}$ represent the predict value and real value of the $i$-step respectively. Table 4 shows the train and test time on the five datasets  [^3] All the experiments were conducted on a server equipped with dual Intel(R) Xeon(R) CPU E5-2678 v3 processors (2.50GHz, 24 cores, 48 threads) and 62 GB of RAM., and Table 3 summarizes our experimental results, which shows the tested results on five datasets using three different initial construction. Here, “origin” represents the original reservoirs initialized by different approaches. A degree-preserving random rewiring baseline (“Random Rewiring” in the Table) was implemented using repeated edge swaps, ensuring identical in-degree and out-degree distributions. The swapping time is set to 0.3 times the number of edges. The spectral radius is re-normalized to 0.8 after rewiring, consistent with our optimization method. As reported in Table 3, the random rewiring fails in many cases, which may make the RMSE become larger, while ours can reduce RMSE in all three initialization methods and five datasets. As for our optimization method, for the Random initialization, the initial reservoir digraphs are sparse and hence show weak memory capacities and prediction abilities at first, even if the dataset has obvious topological features. However, our experimental results demonstrate that even highly sparse reservoirs can achieve substantially improved prediction performance through adding the number of rings within their digraph architecture. And for the small-world and scale-free methods, since in some studies they have been proven to be effective initialization ways of reservoirs [deng2006complex, kawai2019small, cui2012architecture], we find that they have already achieve relative good performance of prediction on some datasets. But for these cases, we have also observed that after adding the number of rings, if the original reservoir already possessed great prediction capability (such as the small-world and scale-free method on the NARMA datasets), the modified results can maintain such good capability, i.e., obtaining similar RMSE or lower RMSE. On the contrary, for reservoirs with initially poor prediction capabilities (such as the Mackey-Glass dataset), the optimization process proves to be even more useful, where the modified reservoirs experience a significant enhancement in their prediction power.
@@
Mackey-Glass 1.8 0.6 1.2 0.6 1.7 0.8
MSO 1.4 0.5 1.7 0.5 1.6 0.6
Lorenz System 1.4 0.7 1.9 0.7 2.0 0.5
NARMA 1.4 0.9 1.8 0.6 1.9 0.8
ETT 1.2 0.5 1.9 0.6 1.8 0.7
Moreover, our optimization method demonstrates varying predictive gains across the evaluated datasets, revealing a compelling Topological Resonance effect that scales with the quantitative Periodicity Index ($PI$) (Fig. 7). The Mackey-Glass dataset, possessing the highest $PI$ (0.1696) and prominent loop features, achieves the most dramatic RMSE reduction (a $\approx 90\%$ improvement for random initialization), suggesting that the GLMY optimization effectively aligns the reservoir’s engineered cycles with the data’s inherent cyclic dynamics. This sensitivity to topological persistence is further corroborated by the results on MSO, Lorenz, and NARMA. These datasets exhibit significantly lower $PI$ values (ranging from 0.0183 to 0.0269) and, as expected, show more constrained performance enhancements. Notably, an interesting observation is found in the ETT dataset. Despite its low $PI$ (0.0135), it exhibits a substantial performance gain (RMSE reduced by $\approx 56\%$ for random initialization). This indicates that our method enhances reservoir performance through two distinct pathways: (1) Topological Resonance: For periodic data, the engineered cycles in the reservoir “resonate” with the data’s cyclic attractor. (2) Memory Extension: As shown in Table 2, the GLMY optimization significantly boosts the Memory Capacity (MC) (from 18.06 to 51.04). Since the ETT task involves long-term forecasting, it benefits more from the extended fading memory provided by the ring-like structures, even in the absence of strong periodicity. This dual-pathway mechanism explains the versatility of our method across diverse datasets.
@@ 5.6 Discussion
5.6.1 Comparison with Traditional Topological and Spectral Approaches
While topological data analysis has been applied to analyze the complexity and decision boundaries of neural networks [rieck2019neural, naitzat2020topology], and spectral graph theory is widely used to ensure the stability of reservoir computing (e.g., the spectral radius condition) [jaeger2001echo, lukovsevivcius2012practical], they exhibit fundamental limitations when applied to the structural optimization of Reservoir Computing (RC). First, traditional persistent homology typically treats networks as undirected graphs or metric spaces. The resulting topological invariants, such as standard Betti numbers ($\beta_{1}$), quantify the number of “holes” or cycles without regard to edge direction. In the context of RC, this is insufficient because an undirected cycle does not necessarily imply a feedback loop, which is the physical mechanism for memory retention. A “hole” in the undirected sense might simply be a feedforward structure, whereas only directed cycles contribute to the recurrent dynamics essential for fading memory. GLMY homology, by contrast, is intrinsically designed for digraphs. It rigorously distinguishes between non-cyclic structures and true directed cycles, ensuring that the optimized topology directly enhances the information recurrence capabilities of the reservoir. Second, spectral graph theory, which analyzes the eigenvalues (spectra) of the Laplacian or Adjacency matrices, provides valuable insights into global properties like network connectivity and stability (e.g., the spectral radius). However, spectral metrics are global statistics; they do not explicitly identify or locate specific local structures that can be surgically modified. For instance, two reservoirs with identical spectral radii can have vastly different cycle distributions and memory capacities. The approach grounded in our GLMY theory provides a perspective that is both local and structural, enabling the accurate identification and modification of particular minimal representative cycles (generators of $H_{1}$) to enhance orthogonality. This degree of detail is unattainable through spectral methods.
5.6.2 Quantitative Interplay between Dataset Topology and Performance
The predictive performance of our GLMY theory-based optimization demonstrates a strong dependency on the intrinsic topological characteristics of the input stream, as quantified by the Periodicity Index ($PI$) in Fig. 7. By comparing the performance gains in Table 3 with these $PI$ values, a clear Topological Resonance effect emerges. Specifically, the Mackey-Glass dataset, which possesses the highest $PI$ (0.1696) and exhibits prominent one-dimensional loops in its embedding space, achieves the most dramatic improvement. In the context of reservoir computing, these $H_{1}$ loops represent recurrent structures essential for capturing temporal dependencies. Our method explicitly strengthens these minimal representative cycles, allowing the reservoir’s optimized digraph to structurally “resonate” with the data’s cyclic attractor. This alignment facilitates more accurate detection and storage of periodic patterns, maximizing the model’s forecasting accuracy.
However, our study further reveals that the method’s effectiveness is not solely restricted to high-$PI$ datasets. A crucial observation is the dual-pathway mechanism of the GLMY theory-based optimization. While datasets like MSO, Lorenz, and NARMA (which have lower $PI$ values (0.0183–0.0269)) show more constrained improvements, the ETT dataset presents a compelling case. Despite having the lowest $PI$ (0.0135), ETT achieves a substantial performance gain. This phenomenon can be comprehensively understood by incorporating the Memory Capacity (MC) analysis from Table 2. Even when the “Topological Resonance” is weak due to indistinct loop features, the GLMY theory-based modification significantly expands the reservoir’s fading memory (increasing ETT’s MC from 18.06 to 51.04). Hence, the effectiveness of our method is the result of a joint influence: (1) Explicit Structural Alignment: directly matching 1-dimensional topological features for periodic data; (2) Functional Memory Extension: universally enhancing the network’s information retention capacity via ring-structured optimization. Only when the method can effectively identify and utilize the topological foundation of the dataset, or significantly boost the requisite memory for the task, can it achieve significant improvements. This provides a generalized framework for understanding how structural optimization interacts with the complex dynamics of time-series data, offering valuable insights for integrating topological analysis with complex system modeling.
5.6.3 Limitations
@@
There are some limitations of our method. First, in this study, we exclude consideration of bidirectional edges (referred to as bigons in the context of GLMY homology), and also, our approach neither captures nor modifies boundary triangles and boundary squares, as they do not constitute generators of GLMY homology, despite being cycles. In dense reservoir digraphs, as the connection density is large, boundary triangles can be numerical, making the number of representative cycles small. But for a sparse reservoir digraph, the boundary cycles are significantly less than the representative cycles; the majority of cycles can still be transformed into rings, thereby enhancing the reservoir’s prediction capability. Hence, our method is more suitable for optimizing sparse reservoirs. And since many rings can share common edges (which means their corresponding fundamental cycle submatrices will share common columns), the increased number of rings is not proportional to improved orthogonality. It is still worthwhile studying more efficient and reasonable methods for improving orthogonality.
@@
Additionally, it is still difficult to give a complete theoretical explanation for why our method can enhance the effect of the reservoir, especially when the reservoir is a large-scale random sparse matrix, because its properties are difficult to study. Moreover, the computational efficiency also becomes a concern for dense reservoirs or large reservoirs, as our method computes GLMY homology and modifies edge directions iteratively, which can be time-intensive. Our future research directions will focus on investigating the theoretical relationship between reservoir topology and computational capability, and developing accelerated algorithms for GLMY homology computation and our algorithm.
@@ 6 Conclusions
In this study, we proposed a novel approach to optimize the structure of the reservoirs using GLMY homology theory. We begin by computing the one-dimensional GLMY homology of the reservoir digraph to obtain the corresponding minimal representative cycles. These cycles are then utilized to strategically modify the digraph structure in a compatible way, specifically increasing the number of rings within the reservoir. Experimental results demonstrate that this topological optimization leads to improved predictive performance across our tested datasets, and the proposed method performs better on datasets with significant 1-dimensional topological features.
@@
Data availability
The raw Electricity Transformer Dataset (ETT) is available at https://github.com/zhouhaoyi/ETDataset/tree/main.